OpenStudy (toasty):

Can someone teach me how to factor these types of problems? x2 - 14x + 40

OpenStudy (anonymous):

figure out some numbers you can multiply to get 40 then figure out which two numbers you can use to subtract to get 14 then figure out which of the two numbers need to end up being negative then put the negative number in the parenthesis with a subtraction sign in it and put the positive number in the parenthesis with a plus sign in it (x-?)(x+?)

OpenStudy (anonymous):

\[(x+a)(x+b) = x^2 + (a+b)x + ab\] a+b = 14 ab = 40 find those numbers through trial and error

OpenStudy (anonymous):

a+b = -14**

OpenStudy (anonymous):

In the example both numbers are negative (-ve * -ve = +ve).